OpenStudy (anonymous):

the measure of angle j is half of its complement plus 12 degrees

OpenStudy (anonymous):

let angle j be x degrees, then complement angle=(90-x) x=((90-x)/2)+12 2x-24=90-x 3x=114 x=38 degrees